Exam #1 Review Problems Chapters 1, 2, 3, 4, 5, and 8 11. Prepare an amortization schedule for a three-year loan of $53,000.Please solve, show work, and give detail explanation
Answer #1
Annual Interest Rate 7.00%
Years 3
Payments Per Year 1
Amount $53,000
Equal Annual Pmt Interest pmt Part of principal repaid Remaining principal
Year
1 =(53000*0.07)*((1+0.07)^3)/((1+0.07)^3-1) 20195.74 =53000.00*7% $3,710.00 $16,485.74 $36,514.26
2 =(36514.26*0.07)*((1+0.07)^2)/((1+0.07)^2-1) 20195.74 =36514.26*7% $2,556.00 $17,639.74 $18,874.52
3 =(18874.52*0.07)*((1+0.07)^1)/((1+0.07)^1-1) 20195.74 =18874.52*7% $1,321.22 $18,874.52 $0.00
$1,321.22 would be interest paid in year 3.
Formulae
(P X R) X [(1+R) ^N] / [(1+R) ^N-1] = Equal annual pmt
Total annual pmt - interest amount = principal pmt repaid
Beginning of year principal - annual principal pmt repaid = balance principal at end of year
P - Principal amount
R - Rate of interest
N - Number of years